# Uranus and Neptune as methane planets: producing icy giants from refractory planetesimals

### v1

@@ 5 Summary and Conclusion
Although the process described here would apply universally to the other giant planets as well, for Jupiter and Saturn the fraction of high-Z material is negligible compared with that of H+He (Helled, 2019), and, as a result, the CH mass fraction expected will be too small to significantly affect the resulting models. Additionally, the interior pressures in these giants would easily dissociate most CH into other hydrocarbons or carbon and hydrogen (Ancilotto et al., 1997; Benedetti et al., 1999; Gao et al., 2010; Sherman et al., 2012). On the other hand, Uranus and Neptune, as well as other similar-sized exoplanets, are massive enough to accrete sufficient hydrogen in order to synthesize ample CH but also just small enough to keep most of this methane from pressure dissociation. In addition, they do not accrete too much hydrogen and helium so that the heavier materials remain a significant fraction of the mass. In addition to the fact that Uranus and Neptune might be methane-rich, there is another interesting outcome of our proposed chemistry. Since large amounts of atmospheric hydrogen are required for methanation, the atmospheric H/He ratio should be significantly sub-solar. While H/He ratio in the bulk is difficult to measure using current remote techniques which only probe the uppermost atmosphere (Atreya et al., 2020), a Uranus mission has recently been recognized as a highest priority target by both NASA and ESA, and is expected to launch sometime during the next decade (Fletcher et al., 2022; Mousis et al., 2022; Girija, 2023; Mandt, 2023). We can expect an entry probe to reveal more high quality data about Uranus’ interior, especially if there is some mixing in the atmosphere, dredging up material from deeper within the planet.
@@ 9 Acknowledgements
We wish to thank Marc Brouwers for useful conversations with regard to giant planet growth and evolution. UM acknowledges support by the Niedersächsisches Vorab in the framework of the research cooperation between Israel and Lower Saxony under grant ZN 3630 and grant by MOST-space. MP acknowledges support by a grant from the Pazy Fund of the Israel Atomic Energy Commission. PB acknowledges support from NASA grant EW-18-2-0060.
@@ Appendix A Are large Kuiper belt objects compatible with comet-like composition?
Based on this aforementioned work (Malamud and Prialnik, 2015), it was concluded that the trend is well produced for a rock/water mass ratio of approximately 3, and a similar conclusion was later echoed in a different study (Bierson and Nimmo, 2019). The exact rock/water mass ratio predicted by such models, depends on what one assumes for the rock specific density (i.e. its grain density). According to the former study (Malamud and Prialnik, 2015), two phases of rocky material were considered. The rock was initially considered to be anhydrous, as in comets (Wooden et al., 2017). Later interaction by the flow of liquid water, which was explicitly followed in the code, transformed the rock and rendered it into the hydrated phase. The specific density of the hydrated phase is typically about 20% lower than the anhydrous phase (Malamud and Prialnik, 2013). The anhydrous rock specific density was a free parameter. Values were considered in the range 3250-3500 kg m. A rock/water mass ratio of 3 matched the Kuiper belt density trend, corresponding to the 3250 kg m value. Slightly changing the value to 3000 or 3500 kg m, only modestly changed these results.
According to the above arguments, for the conclusions to apply, it is merely necessary to justify their choice of rock specific density. Taking the mineralogy assumed in the present work, that is, a binary combination of fayalite and graphite, we can calculate what the specific grain density should be. At zero porosity, as well as low pressure and temperature which would be typical of dwarf-planet sized objects, fayalite and graphite have the specific densities =4390 kg m and =2260 kg m, respectively. Assuming an additive volume law, the density of their combined refractory mixture is given by
where and are the respective mass fractions of fayalite and graphite in the mixture, and . From Equation A1, a 50:50 ratio leads to a density of 3000 kg m, and a 63:37 ratio leads to 3250 kg m. These mix ratios, where organic fractions are in the range 37-50%, are in excellent agreement with those proposed in the Section 2 for outer solar system material. The grain specific density of anhydrous refractories is therefore compatible with previous work (Malamud and Prialnik, 2015), and will lead to the assumed rock-rich intermediate-sized KBOs, having a rock/water mass ratio of around 3.
@@
where =917 kg m is the specific density of water ice, and are the respective mass fractions of the original refractories and water in the dwarf planet, and . The residual porosity in the mantle can be zero, or it can be estimated as some relatively small number. We take 15% as an average value between 0 and 30% discussed above. We now substitute , combining Equations A1 and A2, in order to obtain the value of as a function of the other parameters, leading to
Letting =1854 kg m as determined from observations (Stern et al., 2015), and replacing the refractory fraction with the rock/water mass ratio using the relation , we can use the above relation in order to evaluate as a function of . Figure A1 shows four subplots for different choices of mantle residual , up to 15%.
@@ Appendix B Mission Stardust inability to provide bulk carbon content in comet dust
First, the capture in the aerogel during the flyby occurred at a velocity of around 6 km s. The kinetic energy of a particle impacting aerogel at this speed is dominated by the linear kinetic energy term. For organic grains in particular, this energy typically exceeds their vaporization energy by an order of magnitude. Heating must therefore affect (remove or at least alter) cometary organics to a considerable degree during capture. Additionally, the aerogel itself incorporates organic contaminates originating in its manufacture. Since they are impossible to remove completely, this presents a problem in the sample analysis. Finally, the grains which are collected by Stardust spent an estimated 20 minutes to several hours in the coma prior to collection, which is ample time for an organic grain, or at least a significant fraction of it, to dissociate by interacting with stellar photons.
@@ Appendix C Additional information for models with methane
The data shown in Figures 3 and 5 was used in the main text, depicting some statistical properties related to models where CH was used as a second ice. Here we expand on some additional properties of the random models.

### v2

@@ 5 Broader implications and predictions
Although the process described here would apply universally to other giant planets, for Jupiter and Saturn the fraction of high-Z material is negligible compared with that of H+He (Helled, 2019), and, as a result, the CH4 mass fraction expected will be too small to significantly affect the resulting models. Additionally, the interior pressures in these giants would easily dissociate most CH4 into other hydrocarbons or carbon and hydrogen (Ancilotto et al., 1997; Benedetti et al., 1999; Gao et al., 2010; Sherman et al., 2012). On the other hand, Uranus and Neptune, as well as other similar-sized exoplanets, are massive enough to accrete sufficient hydrogen in order to synthesize ample CH4 but also just small enough to keep most of this methane from pressure dissociation. In addition, they do not accrete too much hydrogen and helium so that the heavier materials remain a significant fraction of the mass.
In addition to the fact that Uranus and Neptune might be methane-rich, there is another interesting outcome of our proposed chemistry. Since large amounts of atmospheric hydrogen are required for methanation, the atmospheric H/He ratio should be significantly sub-solar. We note that the Voyager atmospheric helium mass fraction for Uranus (Conrath et al., 1987) and for Neptune (Conrath et al., 1991) were actually found to be consistent with the primordial solar value. However, these measurements might be uncertain, noting that the Voyager values for the helium mass fraction in Jupiter have been revised upwards by the more recent measurements of Galileo (von Zahn and Hunten, 1996), while the corresponding value for Saturn from Cassini differs by a factor of two (Achterberg and Flasar, 2020). Newer and more precise measurements of the helium mass fraction for Uranus and Neptune would therefore be of great interest. Additionally, it is important to note that the H/He ratio in the bulk of the atmosphere is difficult to measure using current remote techniques which only probe the uppermost atmosphere (Atreya et al., 2020). A Uranus mission has recently been recognized as a highest priority target by both NASA and ESA, and is expected to launch sometime during the next decade (Fletcher et al., 2022; Mousis et al., 2022; Girija, 2023; Mandt, 2023). We can expect an entry probe to reveal more high quality data about Uranus’ interior, especially if there is some mixing in the atmosphere, dredging up material from deeper within the planet.
6 Conclusions
Astrophysical observations suggest that objects in the outer solar system, such as comets and Kuiper belt objects, are refractory-rich. Various pieces of evidence are reviewed (Section 1) which show that the mass ratio of refractory to icy materials in these objects is around 3, or higher. The outer planets Uranus and Neptune are, in contrast, referred to as ice giants, because classic models consider them to be water-rich. Hence, our paper formulates a problem - a tension exists between the composition of the outer planets and the planetesimals they likely accreted during their growth. We explore two possible solutions: either ice-poor planetesimals somehow prompted the formation of ice-rich planets, or the outer planets are actually rock-rich.
We first review the refractory composition of outer solar system planetesimals, and find evidence in the literature that they contain a large mass fraction of refractory organics, between 1/3-1/2 (Section 2). This particular composition is rich in carbon, a fact which is later used to facilitate the first solution.
At the same time, we investigate the detailed composition of Uranus and Neptune (Section 3). We use a computer code that generates hundreds of thousands of random density models for their interiors, commensurate with their observed radius, mass and moment of inertia. We formulate an algorithm that matches these random interior models with a particular composition. We find that planets made of iron, silica, water and hydrogen and helium envelope cannot have a rock/water mass ratio greater than 2, given our algorithm restriction of permitting mixing only between materials of adjacent molecular weights (Section 3.1). Mathematically, mixing iron or silica with a large fraction of hydrogen might help overcome this limit, but it is not clear that this mixture is physically viable (Section 3.2). Adding carbon monoxide as a second ice does not aid in overcoming the aforementioned limit. We find that CO mostly replaces the rock and not the water (Section 3.3). Adding methane as a second ice, however, leads to rock/water mass ratios in the range 0.1-10, because methane can replace water in large amounts. But the ensuing methane mass fraction in the random planet interiors, between 10-50%, is very challenging to explain (Section 3.4).
We conclude that the solution of a rock-rich interior for Uranus and Neptune is possible, but it requires a fine-tuned state of material mixing that might not be feasible. We suggest instead that the abundant organic-rich refractories in outer solar system planetesimals are key to a more natural solution, producing copious amounts of methane through chemical reactions in the hydrogen-rich environment of the growing outer planets (Section 4).
Planetesimals or pebbles are accreted into the proto-planet’s atmosphere and facilitate its growth, while dissolving in the atmosphere and forming compositional gradients based on the material molecular weights (Section 4.1). A simplified chemical model is described to facilitate the conversion of C-rich materials into methane (Section 4.2). We compare the models obtained after this chemical methanation, considering a range of free parameters, to our random interior models with methane, and find an excellent agreement (Section 4.3).
A methane-rich interior therefore seems to present a natural solution to the composition tension problem of the outer planets. We find that Uranus and Neptune are massive enough to accrete sufficient hydrogen, allowing the proposed chemistry, but unlike Jupiter and Saturn, they are not so massive that methane will fully pressure-dissociate in the interior. Additionally, the sum of all high-Z materials inside the major gas giants is dwarfed by their large hydrogen and helium fraction. Uranus, Neptune and similar exoplanets are therefore just in the sweet spot that enables methane-rich planets. Because hydrogen is consumed through chemistry, the bulk He/H ratio is predicted to be larger than solar. We propose this as a priority task for future exploration missions of the outer planets (Section 5).
@@ 10 Acknowledgements
We wish to thank Marc Brouwers for useful conversations with regard to giant planet growth and evolution, Vadim Efimchenko for providing information about his fayalite decomposition laboratory experiments and Eugene Gregoryanz for providing information about his laboratory experiments regarding the methanation of graphite/diamond. UM acknowledges support by the Niedersächsisches Vorab in the framework of the research cooperation between Israel and Lower Saxony under grant ZN 3630 and grant by MOST-space. MP acknowledges support by a grant from the Pazy Fund of the Israel Atomic Energy Commission. PB acknowledges support from NASA grant EW-18-2-0060.
@@ Appendix A Are large Kuiper belt objects compatible with comet-like composition?
Based on this aforementioned work (Malamud and Prialnik, 2015), it was concluded that the trend is well produced for a rock/water mass ratio of approximately 3, and a similar conclusion was later echoed in a different study (Bierson and Nimmo, 2019). The exact rock/water mass ratio predicted by such models, depends on what one assumes for the rock specific density (i.e. its grain density). According to the former study (Malamud and Prialnik, 2015), two phases of rocky material were considered. The rock was initially considered to be anhydrous, as in comets (Wooden et al., 2017). Later interaction by the flow of liquid water, which was explicitly followed in the code, transformed the rock and rendered it into the hydrated phase. The specific density of the hydrated phase is typically about 20% lower than the anhydrous phase (Malamud and Prialnik, 2013). The anhydrous rock specific density was a free parameter. Values were considered in the range 3250-3500 kg m-3. A rock/water mass ratio of 3 matched the Kuiper belt density trend, corresponding to the 3250 kg m-3 value. Slightly changing the value to 3000 or 3500 kg m-3, only modestly changed these results.
According to the above arguments, for the conclusions to apply, it is merely necessary to justify their choice of rock specific density. Taking the mineralogy assumed in the present work, that is, a binary combination of fayalite and graphite, we can calculate what the specific grain density should be. At zero porosity, as well as low pressure and temperature which would be typical of dwarf-planet sized objects, fayalite and graphite have the specific densities =4390 kg m-3 and =2260 kg m-3, respectively. Assuming an additive volume law, the density of their combined refractory mixture is given by
where and are the respective mass fractions of fayalite and graphite in the mixture, and . From Equation A1, a 50:50 ratio leads to a density of 3000 kg m-3, and a 63:37 ratio leads to 3250 kg m-3. These mix ratios, where organic fractions are in the range 37-50%, are in excellent agreement with those proposed in the Section 2 for outer solar system material. The grain specific density of anhydrous refractories is therefore compatible with previous work (Malamud and Prialnik, 2015), and will lead to the assumed rock-rich intermediate-sized KBOs, having a rock/water mass ratio of around 3.
@@
where =917 kg m-3 is the specific density of water ice, and are the respective mass fractions of the original refractories and water in the dwarf planet, and . The residual porosity in the mantle can be zero, or it can be estimated as some relatively small number. We take 15% as an average value between 0 and 30% discussed above. We now substitute , combining Equations A1 and A2, in order to obtain the value of as a function of the other parameters, leading to
Letting =1854 kg m-3 as determined from observations (Stern et al., 2015), and replacing the refractory fraction with the rock/water mass ratio using the relation , we can use the above relation in order to evaluate as a function of . Figure A1 shows four subplots for different choices of mantle residual , up to 15%.
@@ Appendix B Mission Stardust inability to provide bulk carbon content in comet dust
First, the capture in the aerogel during the flyby occurred at a velocity of around 6 km s-1. The kinetic energy of a particle impacting aerogel at this speed is dominated by the linear kinetic energy term. For organic grains in particular, this energy typically exceeds their vaporization energy by an order of magnitude. Heating must therefore affect (remove or at least alter) cometary organics to a considerable degree during capture. Additionally, the aerogel itself incorporates organic contaminates originating in its manufacture. Since they are impossible to remove completely, this presents a problem in the sample analysis. Finally, the grains which are collected by Stardust spent an estimated 20 minutes to several hours in the coma prior to collection, which is ample time for an organic grain, or at least a significant fraction of it, to dissociate by interacting with stellar photons.
@@ Appendix C Additional information for models with methane
The data shown in Figures 3 and 5 was used in the main text, depicting some statistical properties related to models where CH4 was used as a second ice. Here we expand on some additional properties of the random models.